OpenStudy (ammarah):

y=1/2x +2 Graph the equation^

OpenStudy (ammarah):

how would the graph look like?

OpenStudy (anonymous):

i can teach you how to graph it

OpenStudy (anonymous):

Plug in zero for y and solve for x and that is the x intercept. Plug in zero for x and solve for y and that is the y intercept

OpenStudy (anonymous):

this is slope intercept form now specifically y intercept .....  b is the y intercept (or where the line crosses the y axis and m is the slope of the line. (rise/ run)

OpenStudy (ammarah):

show me how to graph

OpenStudy (anonymous):

in y = 1/2x + 2 you need to start the line on the y axis ....and the y intercept is what?

OpenStudy (ammarah):

-2,-1,0,1,2

OpenStudy (anonymous):

to graph this you need to start at the y intercept ammarah ....according to the equation y = (m)x + b b is the intercept y = (1/2)x + 2 is your equation so what is the intercept

OpenStudy (anonymous):

 the y axis will have the y intercept on it and the y instercept is 2 so thats where you will start the line . now the slope of the line is the m or rise/run if your m is a whole number it would be that number over 1 in this case its a fraction 1/2 so go up 1 and over to the right 2 and place another point from that point go up again once and over 2 and place another . to go the opposite direction from the original point you need to go down 1 and to the left 2 and place a point. The dots should have made line.